What would be the current price of a zero-coupon bond with a par value of $1,000, a maturity of 15 years and a yield-to-maturity of 8%? Assume semiannual compounding.
Answer #1

Par value = $1,000

Maturity (n) = 15 years, total periods = 15 * 2 = 30 periods

Yield-to-Maturity (i) = 8%= Semi-annual rate = 8/2 = 4%

Current price = Par value * [1/ (1 + i)n]

Current price = $1,000 * [1 / (1 + 0.04)30]

Current price = $308.32